Based on the function
y=f(x)=1,822.8×(1.4^x)
What would be the predicted transistor count for the years 2006 and 2017? x is the number of years after 1971, so for 1971, x = 0.
The Transistor Count for 2006
Round off the final answer to the nearest whole value.
Transistor Count for 2017
Round off the final answer to the nearest whole value.

Answer by Theo(13342)

the function is y = 1,822.8 * 1.4^x
x is the number of years after 1971,
for 1971, x = 0
for 2006, x = 2006 - 1971 = 35
for 2017, x = 2017 - 1971 = 46
for 2006, y = 1,822.8 * 1.4 ^ 35 = 237,257,674.1
for 2017, x = 1,822.8 * 1.4 ^ 46 = 9, 607,904,134.0
those are your solutions.